# Constraints on the trilinear and quartic Higgs couplings from triple Higgs production at the LHC and beyond

### v1

@@ II Current bounds, unitarity and theoretical motivation
The zeroth partial wave at tree level can be calculated as [^1] We project out the zeroth partial waves from the matrix element computed through FeynArts Kublbeck et al. (1990); Hahn (2001) and FormCalc Hahn and Schappacher (2002), which is in agreement with the result of Ref. Liu et al. (2018).
$\begin{split}a_{ii}^{0}=\frac{3M_{H}^{2}\sqrt{s^{2}-4M_{H}^{2}s}}{32\pi s(s-M_{H}^{2})v^{2}}\left[\kappa_{4}(s-M_{H}^{2})-3\kappa_{3}^{2}M_{H}^{2}\right.\\ \left.+\frac{6\kappa_{3}^{2}M_{H}^{2}(s-M_{H}^{2})}{s-4M_{H}^{2}}\log\left(\frac{s}{M_{H}^{2}}-3\right)\right].\end{split}$ (4)
@@
The unitarity bounds on $\kappa_{4}$ are significantly weaker than the ones on $\kappa_{3}$. This feature can be understood from the Effective Field Theory (EFT) perspective, where effects from higher-dimensional operators to the potential are incorporated as an expansion in terms of inverse powers of a UV-scale $\Lambda$ Boudjema and Chopin (1996) (see also the discussion in Ref. Maltoni et al. (2018)),
@@
Figure 3: Correlation of $\kappa_{3}$ and $\kappa_{4}$ for different values of $M=m_{H}$ and $m_{A}=M_{H^{\pm}}$. The light purple area indicates the range of 2HDM predictions for the correlation between $\kappa_{3}$ and $\kappa_{4}$ for $m_{A},M\in[0.3,10]$ TeV. We show specific choices of $M$ as solid lines with varying $m_{A}$, in order to demonstrate the effect on $\kappa_{3}$ and $\kappa_{4}$. The gray (dark) shaded regions are the areas excluded by tree-level perturbative unitarity according to the analysis of Fig. 1, and the shaded light blue region is defined in the same way as in Fig. 1.
@@ III Triple Higgs production at the HL-LHC
We note that the $4b2\gamma$ and $2b4\tau$ final states only produce a few events at $3$/ab, even at relatively large coupling modifiers $\kappa_{3}\gtrsim 4.5$, $\kappa_{4}\gtrsim 30$ (taking into account K-factors of 1.7 de Florian et al. (2020) and tagging efficiencies of all taus and all-but-one $b$-quarks). It is therefore unlikely that these channels will be statistically significant at the HL-LHC, even though they can be highly relevant for colliders utilising higher energies, as shown in Refs. Fuks et al. (2016); Papaefstathiou and Sakurai (2016); Chen et al. (2016). We therefore will not consider these channels further, and instead focus on the $6b$ and $4b2\tau$ channels.
@@ III.1.1 Event generation and pre-selection
We use Madgraph Alwall et al. (2014); Hirschi and Mattelaer (2015) for event generation and modify the provided SM model file in the UFO Degrande et al. (2012) format to introduce the modifications of the trilinear and quartic Higgs couplings $\kappa_{3}$ and $\kappa_{4}$, respectively. [^5] We checked that our (loop-induced) leading order cross sections for the production of three (undecayed) Higgs bosons is in good agreement with Refs. Plehn and Rauch (2005); Binoth et al. (2006). Signal events are generated for $pp\rightarrow hhh$ and are subsequently decayed on-shell with Madspin Artoisenet et al. (2013) in order to obtain the cross section rates. Due to the complexity of the multi-particle final states we generate events with a minimum transverse momentum for the $b$-quarks of $p_{T}(b)>28$ GeV and within the pseudorapidity region |$\absolutevalue{\eta}<2.5$, while we will later impose stricter cuts during the analysis. Additionally, since the signal consists of three on-shell Higgs bosons, we impose a cut on the invariant mass of the process of $\sqrt{\hat{s}}>350$ GeV at generation level.
@@ III.1.2 Graph Embedding and Neural Network Architecture
GNNs, stemming from the idea that certain types of data can be efficiently represented as graphs, have been increasingly utilised in particle physics. Various works have indicated their applicability for BSM-relevant tasks such as event classification Blance and Spannowsky (2020), jet-tagging Qu and Gouskos (2020); Dreyer and Qu (2021), particle reconstruction Pata et al. (2021), identifying anomalies in data arising from BSM interactions Atkinson et al. (2022a); Atkinson et al. (2021) and obtaining constraints on parameters in SMEFT or the $\kappa$-framework Atkinson et al. (2022b); Anisha et al. (2022). [^8] For a detailed list of references in particle physics using GNNs, see Ref. Feickert and Nachman (2021). The latter is what we aim to achieve by performing a fit within the $\kappa_{3}$$\kappa_{4}$ plane after the efficient selection of a signal region from the GNN. Similar architectures using graphs have also been recently utilised by experiments at the LHC, see e.g. Ref. Aad et al. (2023c).
@@ III.1.3 Training and comparison of graph embeddings
We use the EdgeConv implementation from the Deep Graph Library Wang et al. (2020) with PyTorch Paszke et al. (2019) as backend. The graph embedding relies on PyLHE Heinrich et al. to extract events from the Les Houches Events (LHE) files Alwall et al. (2007). In order to compare the different graph embeddings, we use functionality from scikit-learn Pedregosa et al. (2011) to calculate the true and false positive rates at different thresholds [^11] As we perform multi-class classification for the $3b2\tau$ analysis, we binarise the output of the network for the purpose of this comparison., and we show the corresponding Receiver Operating Characterestic (ROC) curves for both channels in Fig. 4.
@@ III.1.4 HL-LHC Results
For simplicity we will use the signal efficiency of the network for the $(\kappa_{3},\kappa_{4})=(1,1)$ point and assume that it will be mostly the same irrespective of the coupling modifier values. This is not unreasonable as a change of $\kappa_{3}$, $\kappa_{4}$ mostly affects the overall cross section which is not used as input for the GNN. For the $5b$ analysis we optimise the signal selection to reduce the false positive rate to $\sim 0.6\%$. In the $3b2\tau$ channel we require the following conditions to be satisfied on the background network scores:
@@
Overall, we observe that the $3b2\tau$ analysis is more sensitive than the $5b$ analysis, and the latter will additionally suffer from further subdominant electroweak contributions to the background that have not been included. However both channels should be utilised in combinations to maximise the significance. Assuming for simplicity zero correlations between the channels and using the Stouffer method Stouffer et al. (1949), we combine the significances as $Z_{\text{comb}}=(Z_{5b}+Z_{3b2\tau})/\sqrt{2}$, giving rise to the contours shown in Fig. 6. While the projected bounds of about ±$\pm 20$ times the predicted value for the quartic Higgs self-coupling in the SM may appear to be quite weak, in view of our discussion above we emphasise that such bounds go much beyond the existing theoretical bounds. Furthermore, deviations of this size in $\kappa_{4}$ are well compatible with the existing experimental bounds on $\kappa_{3}$ according to the correlations between $\kappa_{3}$ and $\kappa_{4}$ that are present in the BSM scenarios analysed above. Regarding the sensitivity to $\kappa_{3}$ from triple Higgs boson production at the HL-LHC, Fig. 6 shows that the expected sensitivity in this channel at the HL-LHC is weaker than the present experimental limits that have been derived from di-Higgs production. Combining this independent set of experimental information on $\kappa_{3}$ with the experimental results from di-Higgs production may nevertheless turn out to be useful. While our analysis may be optimistic in some respects (e.g. we neglect fake taus), on the other hand we note that further developments of the triggers, tagging and reconstruction algorithms of final states could result in higher efficiencies than the values that we have adopted in our analysis, enhancing the significance. The ability to discriminate between jet flavours is highly important for $HHH$ studies (as well as $HH$ studies) and could also allow experiments to study fully hadronic final states where $H$ decays to $W$ bosons.
[…]
$tt(H\rightarrow\tau\tau)$ $3.8$ $0.17$ $0.011$
$WWbbbb$ $31$ $4.6$ $8.1\times 10^{-3}$
$tt(H\rightarrow bb)$ $3.5$ $0.89$ $3.8\times 10^{-3}$
$Zbbbb$ $4.3$ $0.45$ $3.3\times 10^{-4}$
$tt(Z\rightarrow bb)$ $0.77$ $0.15$ $3.1\times 10^{-4}$
$tt(Z\rightarrow\tau\tau)$ $4.7$ $0.080$ $2.2\times 10^{-4}$
$tttt$ $0.38$ $0.091$ $2.1\times 10^{-4}$
Table 1: Background contributions included in the $3b2\tau$ analysis and reduction of the generated cross sections (labelled as “gen.”) after pre-selection cuts (“sel.”) and GNN selection (“NN”).
Figure 6: Projected contours indicating the $1\sigma$ and $2\sigma$ bounds in the $\kappa_{3}$$\kappa_{4}$ plane obtained from a combination of the $5b$ and $3b2\tau$ channels under the assumption that there are no correlations.
@@ III.2 Interpretability of NN scores
The attributions obtained from IGs allow us to interpret the results of the network in terms of the input parameters for each node, as shown in Fig. 7, although some care is necessary when interpreting such results. Quite intuitively the transverse momenta and the energy of the b-jets are relevant parameters that receive high attributions. This is expected since restricting to higher values of $p_{T}$ can help in the discrimination between signal and background (this was also the reason for applying a pre-selection momentum cut). Angular momenta are not so helpful for discrimination; this is not unexpected as we are dealing with scalars. The network additionally utilises the PID of the tau leptons more than the identification of the $b$ quarks; this is likely due to the fact that the di-tau state is correlated with the highly discriminative reconstructed invariant mass of the Higgs boson. We clearly see that the introduction of the reconstructed masses sigificantly boosts the performance of the network, being the most important observable for the signal events. We note that as a reconstructed particle, the Higgs node has been assigned a PID of zero which as required by the ‘dummy’ axiom [^13] The ‘dummy’ axiom states that a variable that is not contributing to the output of the network should have no attribution, ensuring that the attribution is insensitive to irrelevant inputs. It is a standard axiom imposed by interpretation methods (see e.g. Refs. Sundararajan et al. (2017); Shapley (1951)). has no attribution and thus zero contribution to the network results.
@@ IV Reach assessment for lepton colliders and comparison with the HL-LHC
Figure 11: Comparison of the projected 95% CL contours for the $5b$ and $3b2\tau$ analyses at the HL-LHC as well as their combination with the projected 95% CL sensitivities at lepton colliders with different energies (indicated by the different coloured regions). The shaded gray area indicates the region that is excluded by the bound from tree-level perturbative unitarity.
The leptonic collisions deliver clean signals avoiding the large background contamination from QCD that is present at hadron colliders. We checked the backgrounds of the signature $(5b+\text{other})$ from SM processes and found that the cross sections of these background processes are small. Assuming that the selection of the signal region will enforce $p_{T}(b)>30$ GeV, the requirement that one di-bottom pair should be compatible with the mass of the observed Higgs boson and a cut ensuring that the total invariant mass of the final state particles produced in the process is at least $350$ GeV would result in no remaining background events (even with more relaxed cuts, the number of events is negligible when taking $b$-tagging efficiencies into account). We therefore turn to a Poissonian analysis as described in Ref. Workman et al. (2022), where $n$ corresponds to the number of events expected from the SM, i.e. for $(\kappa_{3},\kappa_{4})=(1,1)$. Upper limits on the mean value of the Poisson distribution $\mu$ are then calculated with
@@
The lepton collider projections are compared with our results for the HL-LHC in Fig. 11. We find that the HL-LHC sensitivity for $\kappa_{4}$ is competitive with the one achievable at a $1$ TeV lepton collider such as the ILC. In particular the comparison shows that for negative $\kappa_{4}$ the HL-LHC is expected to have a better sensitivity than a $1$ TeV lepton collider, while a $1$ TeV lepton collider has a higher sensitivity in the large and positive $\kappa_{3}$ and $\kappa_{4}$ region.

### v3

@@ II Current bounds, unitarity and theoretical motivation
The zeroth partial wave at tree level can be calculated as [^1] We project out the zeroth partial waves from the matrix element computed through FeynArts Kublbeck et al. (1990); Hahn (2001) and FormCalc Hahn and Schappacher (2002), which is in agreement with the result of Ref. Liu et al. (2018) (see also Refs. Di Luzio et al. (2017); Chang and Luty (2020); Falkowski and Rattazzi (2019))
$\begin{split}a_{ii}^{0}=\frac{3M_{H}^{2}\sqrt{s^{2}-4M_{H}^{2}s}}{32\pi s(s-M_{H}^{2})v^{2}}\left[-\kappa_{4}(s-M_{H}^{2})-3\kappa_{3}^{2}M_{H}^{2}\right.\\ \left.+\frac{6\kappa_{3}^{2}M_{H}^{2}(s-M_{H}^{2})}{s-4M_{H}^{2}}\log\left(\frac{s}{M_{H}^{2}}-3\right)\right].\end{split}$ (4)
@@
The unitarity bounds on $\kappa_{4}$ are significantly weaker than the ones on $\kappa_{3}$. This feature can be understood from the Effective Field Theory (EFT) perspective, where effects from higher-dimensional operators to the potential are incorporated as an expansion in terms of inverse powers of a UV-scale $\Lambda$ Boudjema and Chopin (1996) (see also the discussion in Refs. Maltoni et al. (2018); Falkowski and Rattazzi (2019))
@@
Figure 3: Correlation of $\kappa_{3}$ and $\kappa_{4}$ for different values of $M=m_{H}$ and $m_{A}=M_{H^{\pm}}$. The light purple area indicates the range of 2HDM one-loop predictions for the correlation between $\kappa_{3}$ and $\kappa_{4}$ for $m_{A},M\in[0.3,10]$ TeV. We show specific choices of $M$ as solid lines with varying $m_{A}$, in order to demonstrate the effect on $\kappa_{3}$ and $\kappa_{4}$. The gray (dark) shaded regions are the areas excluded by tree-level perturbative unitarity according to the analysis of Fig. 1, and the shaded light blue region is defined in the same way as in Fig. 1.
@@ III Triple Higgs production at the HL-LHC
We note that the $4b2\gamma$ and $2b4\tau$ final states only produce a few events at $3$/ab, even at relatively large coupling modifiers $\kappa_{3}\gtrsim 4.5$, $\kappa_{4}\gtrsim 30$ (taking into account K-factors of 1.7 de Florian et al. (2020) and tagging efficiencies of all taus and all-but-one $b$-quarks). It is therefore unlikely that these channels will be statistically significant at the HL-LHC, even though they can be highly relevant for colliders utilising higher energies, as shown in Refs. Fuks et al. (2016); Papaefstathiou and Sakurai (2016); Chen et al. (2016); Bizoń et al. (2019) [^4] For further studies of other channels at higher energies see also Refs. Kilian et al. (2017, 2020). . We therefore will not consider these channels further, and instead focus on the $6b$ and $4b2\tau$ channels.
@@ III.1.1 Event generation and pre-selection
We use Madgraph5_MC@NLO Alwall et al. (2014); Hirschi and Mattelaer (2015) for event generation and modify the provided SM model file in the UFO Degrande et al. (2012) format to introduce the modifications of the trilinear and quartic Higgs couplings $\kappa_{3}$ and $\kappa_{4}$, respectively. [^6] We checked that our (loop-induced) leading order cross sections for the production of three (undecayed) Higgs bosons is in good agreement with Refs. Plehn and Rauch (2005); Binoth et al. (2006). Signal events are generated for $pp\rightarrow hhh$ and are subsequently decayed on-shell with Madspin Artoisenet et al. (2013) in order to obtain the cross section rates. Due to the complexity of the multi-particle final states we generate events with a minimum transverse momentum for the $b$-quarks of $p_{T}(b)>28$ GeV and within the pseudorapidity region |$\absolutevalue{\eta}<2.5$, while we will later impose stricter cuts during the analysis. Additionally, since the signal consists of three on-shell Higgs bosons, we impose a cut on the invariant mass of the process of $\sqrt{\hat{s}}>350$ GeV at generation level.
@@ III.1.2 Graph Embedding and Neural Network Architecture
GNNs, stemming from the idea that certain types of data can be efficiently represented as graphs, have been increasingly utilised in particle physics. Various works have indicated their applicability for BSM-relevant tasks such as event classification Blance and Spannowsky (2020); Abdughani et al. (2021), jet-tagging Qu and Gouskos (2020); Dreyer and Qu (2021), particle reconstruction Pata et al. (2021), identifying anomalies in data arising from BSM interactions Atkinson et al. (2022a); Atkinson et al. (2021) and obtaining constraints on parameters in SMEFT or the $\kappa$-framework Atkinson et al. (2022b); Anisha et al. (2022). [^9] For a detailed list of references in particle physics using GNNs, see Ref. Feickert and Nachman (2021). The latter is what we aim to achieve by performing a fit within the $\kappa_{3}$$\kappa_{4}$ plane after the efficient selection of a signal region from the GNN. Similar architectures using graphs have also been recently utilised by experiments at the LHC, see e.g. Ref. Aad et al. (2023c).
@@ III.1.3 Training and comparison of graph embeddings
We use the EdgeConv implementation from the Deep Graph Library Wang et al. (2020) with PyTorch Paszke et al. (2019) as backend. The graph embedding relies on PyLHE Heinrich et al. to extract events from the Les Houches Events (LHE) files Alwall et al. (2007). In order to compare the different graph embeddings, we use functionality from scikit-learn Pedregosa et al. (2011) to calculate the true and false positive rates at different thresholds [^12] As we perform multi-class classification for the $3b2\tau$ analysis, we binarise the output of the network for the purpose of this comparison., and we show the corresponding Receiver Operating Characteristic (ROC) curves for both channels in Fig. 4.
@@ III.1.4 HL-LHC Results
For simplicity we will use the signal efficiency of the network for the $(\kappa_{3},\kappa_{4})=(1,1)$ point and assume that it will be mostly the same irrespective of the coupling modifier values as our analysis is largely dependent on the cross section rates. Ideally one could train and optimise a network on each point, or alternatively train on event samples from topologies that depend differently on $\kappa_{3}$ and $\kappa_{4}$. [^13] As a simple test we also trained on a sample that includes two signal classes for $(\kappa_{3},\kappa_{4})=(1,0)$ and $(0,1)$ for the $3b2\tau$ analysis (effectively separating diagrams with $\kappa_{3}$ and $\kappa_{4}$ insertions) and tested this setup on different points. While there is an improvement in efficiency for certain points (up to ${}10\%$), for most cases of $(\kappa_{3},\kappa_{4})$ the efficiency is closer to the simpler setup with one signal class. We chose to utilise the latter, since it also enables an easier interpretation of our results in Sec. III.2. For the $5b$ analysis we optimise the signal selection to reduce the false positive rate to $\sim 0.6\%$. In the $3b2\tau$ channel we require the following conditions to be satisfied on the background network scores:
@@
[…]
$tt(h\rightarrow\tau\tau)$ $3.3$ $0.14$ $0.011$
$wwbbbb$ $27$ $4.0$ $7.1\times 10^{-3}$
$tt(h\rightarrow bb)$ $3.0$ $0.78$ $3.3\times 10^{-3}$
$zbbbb$ $3.8$ $0.40$ $2.9\times 10^{-4}$
$tt(z\rightarrow bb)$ $0.67$ $0.13$ $2.7\times 10^{-4}$
$tttt$ $0.33$ $0.080$ $1.8\times 10^{-4}$
$tt(z\rightarrow\tau\tau)$ $4.1$ $0.073$ $9.0\times 10^{-5}$
Table 1: Background contributions included in the $3b2\tau$ analysis and reduction of the generated cross sections (labelled as “gen.”) after pre-selection cuts (“sel.”) and GNN selection (“NN”). $W$-bosons arising from tops are allowed to decay hadronically and $c$-jets can be mis-tagged as $b$-jets with a probability of $0.2$.
@@
Overall, we observe that the $3b2\tau$ analysis is more sensitive than the $5b$ analysis, and the latter will additionally suffer from further subdominant electroweak contributions to the background that have not been included. [^14] We have tested our trained network on a sample of $Zb\bar{b}b\bar{b}$ events and note that the resulting cross section of this background in the signal region is $\sim 3\%$ of the QCD background. We do not include triple boson backgrounds (e.g. $ZZZ$, $HZZ$ and $HZZ$) which have comparable cross sections to the SM $HHH$ process Hirschi and Mattelaer (2015). Using the conservative assumption that such backgrounds will have an efficiency similar to the signal implies ${\cal{O}}(1)$ events which would be negligible compared to our included backgrounds.
Figure 6: The left plot shows the projected contours indicating the $1\sigma$ and $2\sigma$ bounds in the $\kappa_{3}$$\kappa_{4}$ plane obtained from a combination of the $5b$ and $3b2\tau$ channels under the assumption that there are no correlations. The right plot shows the corresponding result where the backgrounds for both channels are increased by 50%.
However both channels should be utilised in combinations to maximise the significance. Assuming for simplicity zero correlations between the channels, we combine the significances as $Z_{\text{comb}}=\sqrt{Z_{5b}^{2}+Z_{3b2\tau}^{2}}$, giving rise to the contours shown in Fig. 6 (left). While the projected bounds of about ±$\pm 20$ times the predicted value for the quartic Higgs self-coupling in the SM may appear to be quite weak, in view of our discussion above we emphasise that such bounds go much beyond the existing theoretical bounds. Furthermore, deviations of this size in $\kappa_{4}$ are well compatible with the existing experimental bounds on $\kappa_{3}$ according to the correlations between $\kappa_{3}$ and $\kappa_{4}$ that are present in the BSM scenarios analysed above. Regarding the sensitivity to $\kappa_{3}$ from triple Higgs boson production at the HL-LHC, Fig. 6 shows that the expected sensitivity in this channel at the HL-LHC is weaker than the present experimental limits that have been derived from di-Higgs production. Combining this independent set of experimental information on $\kappa_{3}$ with the experimental results from di-Higgs production may nevertheless turn out to be useful. While our analysis may be optimistic in some respects (e.g. we neglect fake taus), on the other hand we note that further developments of the triggers, tagging and reconstruction algorithms of final states could result in higher efficiencies than the values that we have adopted in our analysis, enhancing the significance. The ability to discriminate between jet flavours is highly important for $HHH$ studies (as well as $HH$ studies) and could also allow experiments to study fully hadronic final states where $H$ decays to $W$ bosons. On the other hand, we note that even in the case that the backgrounds are increased by $50\%$, the resulting constraints on $\kappa_{3}$ and $\kappa_{4}$ degrade only slightly, as shown in Fig. 6 (right).
@@ III.2 Interpretability of NN scores
The attributions obtained from IGs allow us to interpret the results of the network in terms of the input parameters for each node, as shown in Fig. 7, although some care is necessary when interpreting such results. Quite intuitively the transverse momenta and the energy of the b-jets are relevant parameters that receive high attributions. This is expected since restricting to higher values of $p_{T}$ can help in the discrimination between signal and background (this was also the reason for applying a pre-selection momentum cut). Angular momenta are not so helpful for discrimination; this is not unexpected as we are dealing with scalars. The network additionally utilises the PID of the tau leptons more than the identification of the $b$ quarks; this is likely due to the fact that the di-tau state is correlated with the highly discriminative reconstructed invariant mass of the Higgs boson. We clearly see that the introduction of the reconstructed masses significantly boosts the performance of the network, being the most important observable for the signal events. We note that as a reconstructed particle, the Higgs node has been assigned a PID of zero which as required by the ‘dummy’ axiom [^16] The ‘dummy’ axiom states that a variable that is not contributing to the output of the network should have no attribution, ensuring that the attribution is insensitive to irrelevant inputs. It is a standard axiom imposed by interpretation methods (see e.g. Refs. Sundararajan et al. (2017); Shapley (1951)). has no attribution and thus zero contribution to the network results.
@@ IV Reach assessment for lepton colliders and comparison with the HL-LHC
Figure 11: Comparison of the projected 95% CL contours for the $5b$ and $3b2\tau$ analyses at the HL-LHC with the projected 95% CL sensitivities at lepton colliders with different energies (indicated by the different coloured regions). The shaded gray area indicates the region that is excluded by the bound from tree-level perturbative unitarity.
The leptonic collisions deliver clean signals avoiding the large background contamination from QCD that is present at hadron colliders. We checked the leading order QCD background of the signature $(5b+\text{other})$ and found that the cross sections of these background processes are small. Assuming that the selection of the signal region will enforce $p_{T}(b)>30$ GeV, the requirement that one di-bottom pair should be compatible with the mass of the observed Higgs boson and a cut ensuring that the total invariant mass of the final state particles produced in the process is at least $350$ GeV would result in no remaining background events (even with more relaxed cuts, the number of events is negligible when taking $b$-tagging efficiencies into account). However, similar to Refs. Maltoni et al. (2018); Gonzalez-Lopez et al. (2021); Chiesa et al. (2020) we do not take into account electroweak backgrounds which could be dominant and deserve a dedicated study. We turn to a Poissonian analysis as described in Ref. Workman et al. (2022), where $n$ corresponds to the number of events expected from the SM, i.e. for $(\kappa_{3},\kappa_{4})=(1,1)$. Upper limits on the mean value of the Poisson distribution $\mu$ are then calculated with
@@
The lepton collider projections are compared with our results for the HL-LHC in Fig. 11. We find that the HL-LHC sensitivity for $\kappa_{4}$ is competitive with the one achievable at a $1$ TeV lepton collider such as the ILC. In particular the comparison shows that for negative $\kappa_{4}$ the HL-LHC is expected to have a better sensitivity than a $1$ TeV lepton collider.